# User Simulation for Evaluating Information Access Systems

### v1

@@ 6.4 Simulating Clicks
Note that here we make the assumption that the user has already examined the snippet $R_{i}$. That is, the clicking policy would only be used in simulating a user when the simulated scanning strategy has predicted examination of the result $R_{i}$. Thus, there are multiple scenarios of interaction with each search result in general as shown in Table 6.2.
@@ 6.4.1 Modeling Clicks in Recommender Systems
Much work has been done recently on modeling clicks and predicting clickthrough rates (CTRs) in the context of recommender systems [Zhou:2018:SIGKDD, Zhang:2022:SIGIR] and online advertisement [Yang:2022:IPM]. In this line of work, the goal is to predict how likely a user would click on a recommended item or an advertised item based on both information about the user and information about the items. The intended application is to use the CTR prediction model to optimize the ranking of items for recommendation or advertising instead of simulating user behaviour, but these two goals are essentially the same in the sense that an item predicted to have high CTRs should be recommended to the user. That is, an effective ranking algorithm for recommendation or advertisement applications needs to be able to predict well whether a user would click on an item if it was recommended to them. Because of the emphasis on the use of such predictive models as recommendation algorithms, these models generally will attempt to use as many relevant features as input to the prediction as possible and are often based on deep neural networks. Those models are commonly used to score an item (with or without a user) and rank all the items with the goal of ranking the items with high CTRs on the top. The performance of those models is typically measured using a ranking-based measure such as Area Under Curve (AUC). Popular models include the Deep Interest Network [Zhou:2018:SIGKDD], Wide and Deep Network [Cheng:2016:Workshop], and Deep Factorization Machines [Guo:2017:IJCAI]. These models often attempt to use historical context information of users, which is typically available in the context of recommender system applications, to improve prediction accuracy [Zhang:2022:SIGIR]. Compared with the models reviewed in [Chuklin:2015:Book], these newer models use richer user information and more sophisticated supervised learning models. The review [Yang:2022:IPM] provides an excellent detailed review and comparison of the major models in this direction.
@@ 6.4.2 On Using Click Models in User Simulators
From the perspective of the MDP framework the models for scanning behaviour and clicking actions together can be regarded as defining the policy used by the simulated user agent when interacting with search results. However, to completely plug in the scanning and clicking models into the MDP framework, we need to further study how to model the user state and the search result context in detail as well as to model a user’s decision process when deciding whether to click on a document with consideration of alternatives if the user would not click on a result. For example, a user may also reformulate a query as an alternative way to continue exploring new documents instead of continuing to click on the current results. The decision at least partially depends on a user’s knowledge about the search topic; if the user is an expert, it would be easier to reformulate a query and thus the user may have the tendency of clicking on fewer results than a novice user who might perceive the cost of reformulating a query as higher than simply continuing the examination of additional results. Thus, in general, modeling clicks needs to consider a user’s cognitive state.
@@ 6.5.2 Simulating Document Utility
Most existing work assumes the presence of relevance assessments for query-document pairs, e.g., by utilizing TREC corpora. This can be understood as assuming such information is included in the state variable of the MDP framework when simulating a user. This is appropriate for simulating scenarios where it reasonable to assume that the user has already had excellent knowledge about what constitutes relevant content for the given query, e.g., an expert user or a known item search task.
@@ 6.5.3 Search Goals and Context
A user’s search goal and context may significantly affect how a user processes results. The underlying user goal in the majority of studies is to find as much relevant material as possible withing a predefined amount of time [Paakkonen:2015:CLEF, Baskaya:2013:CIKM, Verberne:2015:ECIR, Maxwell:2016:CIKM] or until some other stopping condition is met [Carterette:2015:ICTIR, Maxwell:2015:CIKM]. Progress is then measured in terms of the number of relevant documents found or accumulated information gain. Even though dynamic stopping decisions have been considered [Maxwell:2019:PhDThesis], the notion of having found “enough” relevant material has so far been captured only in a simplistic manner, by setting a threshold on the desired number of relevant results. Ultimately, users want answers, not documents, therefore modeling satisfaction should extend beyond the unit of documents, for example, by considering key information nuggets. Further, what constitutes a satisfactory answer might depend on the user’s context. Consider the example of a user searching for the “fastest way to get to the airport.” If the person is performing the search at home, with time to plan their journey and access to a desktop computer or laptop, a satisfactory answer would likely include detailed information about different transportation options, such as driving, public transportation, or ride-sharing services. It may also include information on the estimated time it would take to reach the airport using each option, as well as any potential delays or traffic issues to be aware of. However, if the person is already on the move, using a mobile device, and needs an immediate answer, a satisfactory answer may be more focused on quick and easy directions, such as a map with the fastest route highlighted or step-by-step instructions. To date, there has been limited work on considering different search contexts, even though it has been shown that interactions on the SERP are influenced not only by the presence of heterogeneous content and layout, but also by task complexity [Roy:2022:SIGIR]. For example, when facing complex tasks, users tend to interact more with various components of the SERP [Arguello:2012:SIGIR], examine the result list deeper and faster [Thomas:2013:AIRS]. Wu:2020:CIKM find that the presence of direct answers affects users’ interaction patterns differently for factoid vs. complex question types, resulting in fewer clicks for factoid questions.
@@ 6.6 Simulating Stopping Behaviour
During the search process, users make decisions regarding when to stop examining a result list (query-level stopping) and when to stop issuing queries (session-level stopping), cf. Section 6.1. The difference between the two lies in the granularity at which the information seeking process that would be stopped. Specifically, query-level stopping is an action taken by a user during the sub-process of viewing results from a particular query. Session-level stopping may be regarded as a stopping action taken by a user during the higher-level process of a whole information seeking session. Conceptually, stopping at the session level may be regarded as an implicit action taken by the user which triggers the transition of the (mental) state from “during-information-seeking” to “information-need-satisfied” or “abandon.” The modeling of user decisions around stopping may be complex. For example, imagine a user that has made a query-level stopping decision, but their information need has not been satisfied. In such a case, the user may attempt to formulate a new query, but it they are unable to come up with an appropriate query then may decide to stop the whole session. In general, from the perspective of an MDP, stopping can be defined as a special action to stop a current information seeking process (e.g., examining a list). Once the current process is stopped, the state would transition to a different state, in which the user then faces additional actions to take, including again whether to continue or stop the whole process. A state may be characterized by multiple variables denoting different granularity of process and a stopping action can be applied to any of such processes. A consecutive sequence of stoppings may be possible if they are applied to processes of different granularity. Those processes are closely related to tracking task completion and cost accumulation as well as environmental or contextual constraints.
@@ 6.6.1 Stopping Heuristics
The dependency of stopping on the task of a user means that the state variable in the MDP for modeling stopping should also include information regarding (1) the progress toward completion of the task, (2) importance of finishing the task; searching for medical information may be modeled as a more important task than checking out today’s news; due to the higher utility of the former task, the user would be willing to spend more effort on search, thus leading to delayed stopping as compared with the latter case, where a user could easily give up on search, especially if there are environmental constraints, (3) any constraints from the search context or environment (e.g., downloading and viewing a long document on a mobile device might not be possible if the Internet connection is slow or a user does not have enough time to wait for it to be completed), (4) the perceived cost/effort of continuing the browsing or search (e.g., typing a query while driving is more costly (and may not be impossible) but it has lower cost when done at a meeting, whereas using a speech query has lower cost during driving than at a meeting).
@@ 6.6.3 Modeling Other Stopping Decision Points
Note that with more information-dense SERPs, there is also an increased possibility of “good abandonment,” i.e., that the user’s information need is answered by the SERP without needing to click on any of the results [Li:2009:SIGIR, Chuklin:2012:WWW]. Thus, there is a need for more advanced user browsing models that are able to capture such an implicit action/behaviour. Thomas:2018:ADCS present an extended user model that takes into account the full variety of results searchers see, based on the notion of cards, which refer to various SERP elements, including static, clickable, and interactive ones.
6.7 Toolkits and Resources
In the field of recommender systems research, simulation has been extensively used for measuring and the long-term dynamics of recommender ecosystems, and several toolkits have been created for this purpose, including RecLab [Krauth:2020:arXiv], RecoGym [Rohde:2018:arXiv], PyRecGym [Shi:2019:RecSys], RecSim [Ie:2019:arXiv], and RecSim NG [Mladenov:2021:arXiv]. However, these works operate with overly simplistic user models. We refer to [Stavinova:2022:arXiv] for a detailed description and comparison of simulator toolkits for recommender systems.
6.8 Validating Simulators
For the results of simulations to be credible, it needs to be validated that the simulator imitates the behaviour of real users sufficiently well. This is done by comparing simulated users to real humans, using either user studies or historical log data. Existing work varies in terms of how this comparison is made and which user action(s) are considered.
A commonly used strategy is to measure whether a simulated user would lead to similar retrieval performance to what is obtained from real users. For example, for the purpose of evaluating query simulation the retrieval performance of simulated queries can be validated against real queries [Azzopardi:2007:SIGIR, Baskaya:2011:ECIR, Baskaya:2012:SIGIR, Keskustalo:2009:AIRS]. To validate against a more realistic sample of user queries Breuer:2022:ECIR use a dataset comprising of user query variants [Benham:2017:ADCS].
Another strategy is to measure whether a simulated user would produce data that matches the characteristics of real user data, e.g., session statistics [Maxwell:2016:CIKM, Maxwell:2015:CIKM]. For example, Gunther:2021:Sim4IR ask “how such suggestion-based simulated sessions compare to real user sessions in the sense of coherence, realism, and representativeness of the underlying topic.” Smucker:2012:HCIR compare the distribution of summary times (time it takes users to complete various actions) and gain values to those observed from real users.
Finally, it can be measured how well a user simulator can predict data observed in search logs [Carterette:2015:ICTIR, Chuklin:2015:Book].
One challenge is how to interpret the similarity of a simulated user and a real user. How similar must a simulated user be to a real user in order to be useful for evaluation? Azzopardi:2013:SIGIR test the theory proposed in [Azzopardi:2011:SIGIR] and find that the model does not really fit actual user behaviour. Does this mean that the theoretical model is not useful? On the other hand, the work [Paakkonen:2017:IRJ] studies the validity of the common interaction model typically used for simulating multi-query sessions and show that some strategies for simulation appear to be able to predict cumulated gains over a session reasonably well on unseen real user sessions. It is thus possible that sometimes a simulated user may not replicate a real user’s behaviour well but may be sufficient for evaluation purposes.
@@ 6.9 Summary and Future Challenges
Throughout the chapter, we have discussed how the various decisions a user has to make when interacting with search and recommender systems can be modeled in a general MDP framework in terms of actions, states, rewards, and policy. A component user simulation model (e.g., a click model) can be viewed as modeling the actions that a user can potentially take in a particular state of interaction with a system (e.g., a state of interacting with the search results) and the (local) policy that the user may use to decide which action(s) to take in that state (e.g., whether to click a result). Thus, the MDP framework can serve as a powerful and mathematically sound framework for studying user simulation. Different user simulation models and techniques can all be viewed as specific instantiations of the general framework under various assumptions.
@@
Existing work has generally only focused on the explicit high-level states, with some work modeling the cognitive state of users during the process of interaction with a simple list of results. More detailed and broad modeling of the state is needed in the future, which requires (1) modeling a user’s task deeply and refining task workflows; and (2) modeling complex user interfaces, including different layouts as well as conversational interfaces; (3) modeling the context/environment of the search (e.g., time and location); and (4) better understanding of complex processes such as cognition and emotional reactions (such as frustration or impatience).
@@
Policy: The policy of an MDP determines how to choose an action in each state. This is at the core of computationally modeling a user’s behaviour and connects all the other components in the MDP. Given the current state, which provides all the information for deciding an action, including, e.g., a user’s current cognitive and emotional states, the current interface a user is interacting with, the policy would determine the best action for a user to take (which could also be an implicit action that a user takes mentally), a chosen action would then cause the state to transition to a new state and the process repeats in this way. Existing work has modeled the policy in some individual explicit states, such as initial query formulation, scanning of results in a list, clicking on a result, and reformulation of a query. The models of policy can be either simple but interpretable models or machine-learned non-interpretable predictive models of user behaviour.
@@
A major gap is the lack of work on integrating multiple component policies into a unified approach, with a systematic treatment of all the states and their transitions. There have already been some attempts to address this by loosely connecting policy models of the common explicit states in a search system, thus integrating initial and subsequent query formulation/reformulation and click modeling. In the future, tighter integration of component models with full specification of states and state transitions is needed, so as to bring all the research results on individual components together in the MDP framework with the purpose of building more realistic user simulators.
@@
In order to realistically model decision making, it is important to consider how humans tend to employ a high-level strategy to balance costs and benefits, and adopt these trade-offs based on their current context. Such higher-level strategies might be informed by conceptual models of information seeking and decision making (cf. Section 5.1).
@@
We have shown that the MDP framework can be used as a general framework to synthesize both past and future work on simulating user interactions with search and recommender systems. It is worth noting that other mathematical frameworks could also be used, and the exploration of alternatives and the advantages they can offer is an important direction for future research. Nevertheless, the fundamental principles and insights discussed in this chapter, such as the modeling the space of actions and associated costs/rewards for decision making are likely to remain relevant across different modeling frameworks. In particular, as we will further discuss in the next chapter, the MDP framework is also applicable to simulating user interactions with conversational assistants.
Characteristic Description Reference
Patience
Patience to browse through the list of retrieved documents
[Keskustalo:2006:ECIR]
Feedback
Willingness to provide relevance feedback
[Keskustalo:2006:ECIR]
Scanning behaviour
Time allocation between examining summaries and reading documents
[Smucker:2011:HCIR]
Expertise
Search experience
[Paakkonen:2015:CLEF]
Learning speed
Speed of incorporating novel terms into vocabulary
[Camara:2022:ECIR]
Exploration
Willingness to explore different subtopics
[Camara:2022:ECIR]
Tolerance
Willingness to click on a search result snippet
[Camara:2022:ECIR]
Subtopic switching
Strategy employed to navigate through subtopics
[Camara:2022:ECIR]
@@
Finally, we would like to point out the close connection of user simulation and user studies. Many user studies have revealed useful findings about user characteristics and behaviours. In general, this kind of findings have not yet made into formal design of user simulators, though some findings have clearly influenced the design of user simulation models. The topic of user simulation naturally connects user studies and formal model design, therefore we anticipate such integrated work to be critical to eventually develop realistic user simulators. Table 6.3 lists some specific user characteristics that have been considered in past work, but the challenge of incorporating them into policies in a generic manner remains to be solved.

### v2

@@ 6.4 Simulating Clicks
Note that here we make the assumption that the user has already examined the snippet $R_{i}$. That is, the clicking policy would only be used in simulating a user when the simulated scanning strategy has predicted examination of the result $R_{i}$. Thus, there are multiple scenarios of interaction with each search result in general as shown in Table 6.2. Note that the interpretation of viewing a snippet without clicking on the result (to further examine the content of the document) is complicated since there is a possibility that the snippet might have already provided the relevant information that the user is searching for (e.g., a telephone number). In such a case, we cannot assume the snippet to be non-relevant; it is indeed the opposite case. From a user simulation perspective, this adds complexity to the space of actions since there is a non-observable action of “good abandonment” [Li:2009:SIGIR]. With enriched user interfaces, where relevant information can be easily shown to users directly, becoming more common, such special unobservable actions may also become more common. This action is more related to examination of document content than clicking on a document (see Section 6.5) and how to incorporate such actions into an evaluation measure may be complicated, but in general, it needs to be modeled with consideration of the overall utility of the system as suggested in [Machmouchi:2017:CIKM]. The framework we presented in Chapter 4 can be used to define such an overall utility based on an entire interaction sequence of user actions.
@@
In general, the parameters of a click model can be estimated based on search log data. When the click model involves only parameters representing observable events, the parameters can be easily estimated by normalizing the counts of the corresponding events of those parameters based on the analytical form of the Maximum Likelihood (ML) estimator. When a model involves a latent variable that is not observable, which are clearly needed to model the scanning behaviour of a user, there is generally no analytical solution to the ML-estimation problem, and an iterative numerical algorithm such as the EM algorithm is often used to compute the ML estimate and estimate all the parameter values [Chuklin:2015:Book].
@@ 6.4.1 Modeling Clicks in Recommender Systems
Much work has been done recently on modeling clicks and predicting clickthrough rates (CTRs) in the context of recommender systems [Zhou:2018:SIGKDD, Zhang:2022:SIGIR] and online advertisement [Yang:2022:IPM]. In this line of work, the goal is to predict how likely a user would click on a recommended item or an advertised item based on both information about the user and information about the items. The intended application is to use the CTR prediction model to optimize the ranking of items for recommendation or advertising instead of simulating user behaviour, but these two goals are essentially the same in the sense that an item predicted to have high CTRs should be recommended to the user. That is, an effective ranking algorithm for recommendation or advertisement applications needs to be able to predict well whether a user would click on an item if it was recommended to them. In this sense, such models could technically be regarded as modeling user clicks; indeed, they are generally evaluated based on how accurate their prediction of user clicks is. However, because of the emphasis on the use of such predictive models as recommendation algorithms rather than user simulators, these models generally will attempt to use as many relevant features as input to the prediction as possible and are often based on deep neural networks. Those models are commonly used to score an item (with or without a user) and rank all the items with the goal of ranking the items with high CTRs on the top. The performance of those models is typically measured using a ranking-based measure such as Area Under Curve (AUC). Popular models include the Deep Interest Network [Zhou:2018:SIGKDD], Wide and Deep Network [Cheng:2016:Workshop], and Deep Factorization Machines [Guo:2017:IJCAI]. These models often attempt to use historical context information of users, which is typically available in the context of recommender system applications, to improve prediction accuracy [Zhang:2022:SIGIR]. Compared with the models reviewed in [Chuklin:2015:Book], these newer models use richer user information and more sophisticated supervised learning models. The review [Yang:2022:IPM] provides an excellent detailed review and comparison of the major models in this direction.
While most existing work can be regarded as non-interpretable approaches as discussed above, interpretable models have also been recently explored. For example, [Rahdari:2024:TORS] develop a carousel click model for modeling clicking behaviour on carousel-based recommender interfaces, which generalizes the traditional cascade model [Craswell:2008:WSDM]. Another interesting recent line of work on choice models [Hazrati:2022:IPM, Hazrati:2024:UMUAI] has taken a more interpretable approach to simulate clicking in recommender systems. For example, [Hazrati:2022:IPM, Hazrati:2024:UMUAI] use three choice models, Age-CM, Popularity-CM, and Rating-CM, each modeling a somewhat different clicking behaviour. Specifically, Hazrati:2024:UMUAI describe them as: “Age-CM models users that tend to choose newer items among the recommended ones. Popularity-CM instead models users that tend to choose more popular items. Finally, Rating-CM models users that choose items with higher ratings.” A baseline choice model, Base-CM, is also considered where the simulated user would faithfully accept recommended results (thus potentially containing position bias). The authors employ these models in simulation experiments involving a population of simulated users repeatedly interacting with a simulated recommender system. Such a simulation experiment enables the analysis of many questions about the community-wide behaviour of users and systems over a long period with some interesting findings about the complicated interactions between users’ choices and the recommender systems. From evaluation perspective, this line of work demonstrates the feasibility of using user simulation to evaluate the overall utility and impact of a recommender system if deployed to a community of users. Although the perspective of evaluation we have taken in the book is more on comparing two interactive systems, such a broader perspective of user simulation for analysis of a complex interactive ecosystem of users and systems is no doubt a very interesting future research direction, where user simulation also plays a critical role.
@@ 6.4.2 On Using Click Models in User Simulators
From the perspective of the MDP framework, the models for scanning behaviour and clicking actions together can be regarded as defining the policy used by the simulated user agent when interacting with search results. However, to completely plug in the scanning and clicking models into the MDP framework, we need to further study how to model the user state and the search result context in detail as well as to model a user’s decision process when deciding whether to click on a document with consideration of alternatives if the user would not click on a result. For example, a user may also reformulate a query as an alternative way to continue exploring new documents instead of continuing to click on the current results. The decision at least partially depends on a user’s knowledge about the search topic; if the user is an expert, it would be easier to reformulate a query and thus the user may have the tendency of clicking on fewer results than a novice user who might perceive the cost of reformulating a query as higher than simply continuing the examination of additional results. Thus, in general, modeling clicks needs to consider a user’s cognitive state.
@@ 6.5.2 Simulating Document Utility
Most existing work assumes the presence of relevance assessments for query-document pairs, e.g., by utilizing TREC corpora. This can be understood as assuming such information is included in the state variable of the MDP framework when simulating a user. This is appropriate for simulating scenarios where it is reasonable to assume that the user has already had excellent knowledge about what constitutes relevant content for the given query, e.g., an expert user or a known item search task.
@@
It is worth pointing out the connection to recent work on employing large language models for estimating relevance judgments, which may be considered a data-driven and non-interpretable simulation approach [Faggioli:2023:ICTIR]. However, due to the large amount of information “seen" by LLMs during the pre-training stage, they might be more suitable for simulating “expert users” who possess substantial knowledge on the topic, rather than those who are completely new to it (see Section 9.3 for further discussion on the use of LLMs for simulation).
@@ 6.5.3 Search Goals and Context
A user’s search goal and context may significantly affect how a user processes results. The underlying user goal in the majority of studies is to find as much relevant material as possible withing a predefined amount of time [Paakkonen:2015:CLEF, Baskaya:2013:CIKM, Verberne:2015:ECIR, Maxwell:2016:CIKM] or until some other stopping condition is met [Carterette:2015:ICTIR, Maxwell:2015:CIKM]. Progress is then measured in terms of the number of relevant documents found or accumulated information gain. Ultimately, users want answers, not documents, therefore modeling satisfaction should extend beyond the unit of documents, for example, by considering key information nuggets. Further, what constitutes a satisfactory answer might depend on the user’s context. Consider the example of a user searching for the “fastest way to get to the airport.” If the person is performing the search at home, with time to plan their journey and access to a desktop computer or laptop, a satisfactory answer would likely include detailed information about different transportation options, such as driving, public transportation, or ride-sharing services. It may also include information on the estimated time it would take to reach the airport using each option, as well as any potential delays or traffic issues to be aware of. However, if the person is already on the move, using a mobile device, and needs an immediate answer, a satisfactory answer may be more focused on quick and easy directions, such as a map with the fastest route highlighted or step-by-step instructions. To date, there has been limited work on considering different search contexts, even though it has been shown that interactions on the SERP are influenced not only by the presence of heterogeneous content and layout, but also by task complexity [Roy:2022:SIGIR]. For example, when facing complex tasks, users tend to interact more with various components of the SERP [Arguello:2012:SIGIR], examine the result list deeper and faster [Thomas:2013:AIRS]. Wu:2020:CIKM find that the presence of direct answers affects users’ interaction patterns differently for factoid vs. complex question types, resulting in fewer clicks for factoid questions.
@@ 6.6 Simulating Stopping Behaviour
During the search process, users make decisions regarding when to stop examining a result list (query-level stopping) and when to stop issuing queries (session-level stopping), cf. Section 6.1. The difference between the two lies in the granularity at which the information seeking process that would be stopped. Specifically, query-level stopping is an action taken by a user during the sub-process of viewing results from a particular query. Session-level stopping may be regarded as a stopping action taken by a user during the higher-level process of a whole information seeking session. Conceptually, stopping at the session level may be regarded as an implicit action taken by the user which triggers the transition of the (mental) state from “during-information-seeking” to “information-need-satisfied” or “abandon.” The modeling of user decisions around stopping may be complex. For example, imagine a user that has made a query-level stopping decision, but their information need has not been satisfied. In such a case, the user may attempt to formulate a new query, but if they are unable to come up with an appropriate query then they may decide to stop the whole session. In general, from the perspective of an MDP, stopping can be defined as a special action to stop a current information seeking process (e.g., examining a list). Once the current process is stopped, the state would transition to a different state, in which the user then faces additional actions to take, including again whether to continue or stop the whole process. A state may be characterized by multiple variables denoting different granularity of process and a stopping action can be applied to any of such processes. A consecutive sequence of stoppings may be possible if they are applied to processes of different granularity. Those processes are closely related to tracking task completion and cost accumulation as well as environmental or contextual constraints.
@@ 6.6.1 Stopping Heuristics
The dependency of stopping on the task of a user means that the state variable in the MDP for modeling stopping should also include information regarding (1) the progress toward completion of the task, (2) importance of finishing the task; searching for medical information may be modeled as a more important task than checking out today’s news; due to the higher utility of the former task, the user would be willing to spend more effort on search, thus leading to delayed stopping as compared with the latter case, where a user could easily give up on search, especially if there are environmental constraints, (3) any constraints from the search context or environment (e.g., downloading and viewing a long document on a mobile device might not be possible if the Internet connection is slow or a user does not have enough time to wait for it to be completed), (4) the perceived cost/effort of continuing the browsing or search (e.g., typing a query while driving is more costly (and may be impossible) but it has lower cost when done at a meeting, whereas using a speech query has lower cost during driving than at a meeting).
@@ 6.6.2 Modeling Query-level Stopping
Note that query-level stopping may be affected by whether the user can easily reformulate the query; intuitively, if the user can easily come up with a new, potentially more effective query, then they may tend to stop browsing the current search results earlier, but if the user has already reformulated a few queries, they may have no choice but to continue browsing the current list.
@@ 6.6.3 Modeling Other Stopping Decision Points
Note that with more information-dense SERPs, there is also an increased possibility of “good abandonment,” i.e., that the user’s information need is answered by the SERP without needing to click on any of the results [Li:2009:SIGIR, Chuklin:2012:WWW]. Thus, there is a need for more advanced user browsing models that are able to capture such implicit action. Thomas:2018:ADCS present an extended user model that takes into account the full variety of results searchers see, based on the notion of cards, which refer to various SERP elements, including static, clickable, and interactive ones.
@@ 6.7 Summary and Future Challenges
Throughout the chapter, we have discussed how the various decisions a user has to make when interacting with search and recommender systems can be modeled in a general MDP framework in terms of actions, states, rewards, and policy. A component user simulation model (e.g., a click model) can be viewed as modeling the actions that a user can potentially take in a particular state of interaction with a system (e.g., a state of interacting with the search results) and the (local) policy that the user may use to decide which action(s) to take in that state (e.g., whether to click a result). Thus, the MDP framework can serve as a powerful and mathematically sound framework for studying user simulation. Different user simulation models and techniques can all be viewed as specific instantiations of the general framework under various assumptions, and machine learning algorithms, particularly model-based reinforcement learning [Moerland:2023:Survey] and imitation learning [Hussein:2017:CSUR, Labhishetty:2023:thesis], can be naturally used to train an interpretable user simulator based on user data.
@@
Existing work has generally only focused on the explicit high-level states, with some work modeling the cognitive state of users during the process of interaction with a simple list of results. More detailed and broad modeling of the state is needed in the future, which requires (1) modeling a user’s task deeply and refining task workflows; (2) modeling complex user interfaces, including different layouts as well as conversational interfaces; (3) modeling the context/environment of the search (e.g., time and location); and (4) better understanding of complex processes such as cognition and emotional reactions (such as frustration or impatience).
@@
Policy: The policy of an MDP determines how to choose an action in each state. This is at the core of computationally modeling a user’s behaviour and connects all the other components in the MDP. Given the current state, which provides all the information for deciding an action, including, e.g., a user’s current cognitive and emotional states, the current interface a user is interacting with, the policy would determine the best action for a user to take (which could also be an implicit action that a user takes mentally). The chosen action would then cause the state to transition to a new state and the process repeats in this way. Since existing work has mostly focused on simulating some specific actions of a user, the models developed in existing work can only be used to define the policy function for some special states, such as the state in which a user would formulate an initial query, the state in which a user would scan results in a list, or click on a result, and the state in which a user would decide whether to reformulate a query. The models of policy can be either simple but interpretable models or machine-learned non-interpretable predictive models of user behaviour.
@@
A major gap is the lack of work on integrating multiple component policies into a unified approach, with a systematic treatment of all the states and their transitions, thus completely defining the policy function. There have already been some attempts to address this by loosely connecting policy models of the common explicit states in a search system, thus integrating initial and subsequent query formulation/reformulation and click modeling. In the future, tighter integration of component models with full specification of states and state transitions is needed, so as to bring all the research results on individual components together in the MDP framework with the purpose of building more realistic user simulators.
@@
In order to realistically model decision making, it is important to consider how humans tend to employ a high-level strategy to balance costs and benefits, and adopt these trade-offs based on their current context. Such higher-level strategies might be informed by conceptual models of information seeking and decision making (cf. Sections 5.1 and  5.2).
@@
We have shown that the MDP framework can be used as a general framework to synthesize both past and future work on simulating user interactions with search and recommender systems. It is worth noting that other mathematical frameworks could also be used, and the exploration of alternatives and the advantages they can offer is an important direction for future research. Nevertheless, the fundamental principles and insights discussed in this chapter, such as modeling the space of actions and associated costs/rewards for decision making, are likely to remain relevant across different modeling frameworks. In particular, as we will further discuss in the next chapter, the MDP framework is also applicable to simulating user interactions with conversational assistants.
@@
Finally, we would like to point out the close connection of user simulation and user studies. Many user studies have revealed useful findings about user characteristics and behaviours. In general, these findings have not yet made into formal design of user simulators, though some findings have clearly influenced the design of user simulation models. The topic of user simulation naturally connects user studies and formal model design, therefore we anticipate such integrated work to be critical to eventually develop realistic user simulators. Table 8.1 lists some specific user characteristics that have been considered in past work, but the challenge of incorporating them into policies in a generic manner remains to be addressed.